### v1

@@ E.1 Alignment and continuity discussion
Assume that the clean event $G$ occurs. This occurs with probability at least $1-(2\cdot|\mathcal{A}|+1)\cdot T^{-3}$ (Lemma E.4), so the clean event not occurring counts negligibly towards regret.
We apply Lemma E.6 to see that at time step $t$, it holds that for $\epsilon_{t}=\Theta(\sqrt{|\mathcal{B}|\cdot\log T/n_{t}(a_{t})})$, it holds that
@@
where (A) follows from Lemma B.1. This gives the desired upper bound.
@@
Appendix F Proofs for Section 6
F.1 Proof of Theorem 6.2
@@ Appendix F Proofs for Section 6
The proof follows a similar argument to the proof of Theorem 4.4 and borrows some lemmas from Appendix D.2
@@
Assume that the clean event $G$ holds. This occurs with probability at least $1-(|\mathcal{A}|\cdot|\mathcal{B}|+|\mathcal{A}|)T^{-3}$ (Lemma D.1), so the clean event not occuring counts negligibly towards regret.
@@
We focus on $t>E_{2}\cdot|\mathcal{B}|\cdot|\mathcal{A}|+E_{1}\cdot|\mathcal{A}|$ for the remainder of the analysis. Our main ingredient is Lemma D.2. Note that $\epsilon^{*}=\Theta\left(\max\left(\frac{\sqrt{\log T}}{\sqrt{E_{1}}},\frac{\sqrt{\log T}}{\sqrt{E_{2}}}\right)\right)=\Theta\left((|\mathcal{A}|\cdot|\mathcal{B}|\cdot(\log T))^{\eta/2}\cdot(c\cdot T)^{-\eta/2}\right)$ based on the settings of $E_{1}$ and $E_{2}$. The regret of the leader can be bounded as:
@@
where (A) follows from Lemma D.2 and (B) uses the fact that $c\geq 1$ and $d\leq 1$. The regret of the follower can similarly be bounded as:
@@
where (B) follows from Lemma D.2. This proves the desired result.
F.2 Proof of Theorem 6.3
@@
The proof follows a similar argument to the proof of Theorem 4.4 and borrows some lemmas from Appendix D.3
@@
Assume that the clean event $G$ holds. This occurs with probability at least $1-(1+|\mathcal{A}|)T^{-3}$ (Lemma D.3), so the clean event not occurring counts negligibly towards regret.
@@
$\beta_{1}^{\text{tol}}\cdot(T-E\cdot\left|\mathcal{A}\right|)-\sum_{t=E\cdot|\mathcal{A}|}^{T}v^{1}_{a_{t},b_{t}}$
@@
$=\sum_{a\in\mathcal{A}}\sum_{t\in T_{a}}\left(c\cdot(\epsilon^{*})^{d}+\max_{a\in\mathcal{A}}\min_{b\in\mathcal{B}_{\epsilon^{*}}(a)}v_{1}(a,b)-v_{1}(a_{t},b_{t})\right)$
$\leq|\mathcal{A}|+\sum_{a\in\mathcal{A}}\sum_{t\in n_{E\cdot|\mathcal{A}|,T}(a)\setminus\left\{\max(S(a))\right\}}\left(c\cdot(\epsilon^{*})^{d}+\max_{a\in\mathcal{A}}\min_{b\in\mathcal{B}_{\epsilon^{*}}(a)}v_{1}(a,b)-v_{1}(a_{t},b_{t})\right)$
@@
$+\underbrace{\sum_{a\in\mathcal{A}}(n_{E\cdot|\mathcal{A}|,T}(a)-1)\cdot\left(\max_{a\in\mathcal{A}}\min_{b\in\mathcal{B}_{\epsilon^{*}}(a)}v_{1}(a,b)-\frac{1}{n_{E\cdot|\mathcal{A}|,T}(a)-1}\sum_{t\in S(a)\setminus\left\{\max(S(a))\right\}}\left(v_{1}(a_{t},b_{t})\right)\right)}_{(2)}$
@@
$\sum_{a\in\mathcal{A}}(n_{E\cdot|\mathcal{A}|,T}(a)-1)\cdot\left(\max_{a\in\mathcal{A}}\min_{b\in\mathcal{B}_{\epsilon^{*}}(a)}v_{1}(a,b)-\frac{1}{n_{E\cdot|\mathcal{A}|,T}(a)-1}\sum_{t\in S(a)\setminus\left\{\max(S(a))\right\}}\left(v^{1}_{a_{t},b_{t}}\right)\right)$
$\leq\sum_{a\in\mathcal{A}}(n_{E\cdot|\mathcal{A}|,T}(a)-1)\cdot\frac{20\sqrt{\log T}}{\sqrt{n_{E\cdot|\mathcal{A}|,T}(a)}-1}$
$\leq O\left(\sqrt{\log T}\cdot\sum_{a\in\mathcal{A}}\sqrt{n_{E\cdot|\mathcal{A}|,T}(a)-1}\right)$
@@
where the first inequality uses Lemma D.4 and the last inequality uses Jensen’s inequality.
@@
However, by Lemma D.4, we can also lower bound the average reward across all time steps (except for the last time step) where $a^{\prime}$ is pulled in terms of $n_{E\cdot|\mathcal{A}|,T}(a^{\prime})$ as follows:
@@
Appendix G Proofs for Section 7
@@
We use the following notation in the proof. Let $n_{t}(a)$ be the number of times that arm $a$ has been pulled up prior to time step $t$. Following Appendix B, the follower’s history can be represented as:
@@
$\mathbb{P}\left[\forall t\in[T],a\in\mathcal{A}\mid v_{2}(a_{t},b_{t})\geq\max_{b\in\mathcal{B}}v_{2}(a_{t},b)-g(n_{t}(a)+1,T)\right]\geq 1-|\mathcal{A}|\cdot T^{-3},$
@@
Using Lemma G.1, it suffices to analyze the high-probability instantaneous regret and high-probability anytime regret of the following standard bandit algorithms as single-bandit learners with arms $\mathcal{B}$, mean rewards $v_{2}(b)$, and stochastic rewards $r_{2,t}(b)$. In the proofs, we let $n_{t}(b)$ denote the number of times that arm $b$ has been pulled prior to time step $t$.
@@
By Lemma G.1, it suffices to show the bound for ActiveArmElimination (using phase lengths $M_{i}=\Theta(\log(T)\cdot 2^{2i})$) as a single-bandit learner with arms $\mathcal{B}$, mean rewards $v_{2}(b)$, and stochastic rewards $r_{2,t}(b)$. We let $n_{t}(b)$ denote the number of times that arm $b$ has been pulled prior to time step $t$ in the current phase. Let $\hat{v}_{2,t}(b)$ denote the empirical mean reward for arm $b$ over the rewards observed prior to time step $t$ in the previous (last completed) phase. Let $\mathcal{B}^{\prime}_{t,\text{curr}}$ be the set of arms active in the current phase, and let $\mathcal{B}^{\prime}_{t,\text{prev}}$ be the set of arms active in the previous (last completed) phase. For each time step $t$, let $s^{\prime}_{t}$ denote the index of the previous (last completed) phase at time step $t$.
@@
By Lemma G.1, it suffices to show the instantaneous regret bound for ExploreThenCommit as a single-bandit learner with arms $\mathcal{B}$, mean rewards $v_{2}(b)$, and stochastic rewards $r_{2,t}(b)$. We let $n_{t}(b)$ denote the number of times that arm $b$ has been pulled prior to time step $t$. Let $\hat{v}_{2,t}(b)$ denote the empirical mean reward for arm $b$ over the rewards observed prior to time step $t$.
@@
By Lemma G.1, it suffices to show the anytime regret bound for UCB as a single-bandit learner with arms $\mathcal{B}$, mean rewards $v_{2}(b)$, and stochastic rewards $r_{2,t}(b)$. We let $n_{t}(b)$ denote the number of times that arm $b$ has been pulled prior to time step $t$. Let $\hat{v}_{2,t}(b)$ denote the empirical mean reward for arm $b$ over the rewards observed prior to time step $t$.
@@
We condition on $G$ for the remainder of the analysis. Since the arm with highest upper confidence bound is always chosen and since $G$ holds, the selected arm $b_{t}$’s true mean $v_{2}(b_{t})$ falls within the $2\cdot\sqrt{\frac{\log(T)}{n_{t}(b_{t})}}$ bound. By Lemma B.1, this means that the regret at any time step $t$ for any arm $a\in\mathcal{A}$ is upper bounded by:

### v2

@@ D.1 Alignment and continuity discussion
Assume that the clean event $G$ occurs. This occurs with probability at least $1-(2\cdot|\mathcal{A}|+1)\cdot T^{-3}$ (Lemma D.4), so the clean event not occurring counts negligibly towards regret.
We apply Lemma D.6 to see that at time step $t$, it holds that for $\epsilon_{t}=\Theta(\sqrt{|\mathcal{B}|\cdot\log T/n_{t}(a_{t})})$, it holds that
@@
where (A) follows from Lemma A.1. This gives the desired upper bound.
@@
Appendix E Proofs for Section 6
E.1 Proof of Theorem 6.2
@@ Appendix E Proofs for Section 6
The proof follows a similar argument to the proof of Theorem 4.4 and borrows some lemmas from Appendix C.2
@@
Assume that the clean event $G$ holds. This occurs with probability at least $1-(|\mathcal{A}|\cdot|\mathcal{B}|+|\mathcal{A}|)T^{-3}$ (Lemma C.1), so the clean event not occuring counts negligibly towards regret.
@@
We focus on $t>E_{2}\cdot|\mathcal{B}|\cdot|\mathcal{A}|+E_{1}\cdot|\mathcal{A}|$ for the remainder of the analysis. Our main ingredient is Lemma C.2. Note that $\epsilon^{*}=\Theta\left(\max\left(\frac{\sqrt{\log T}}{\sqrt{E_{1}}},\frac{\sqrt{\log T}}{\sqrt{E_{2}}}\right)\right)=\Theta\left((|\mathcal{A}|\cdot|\mathcal{B}|\cdot(\log T))^{\eta/2}\cdot(c\cdot T)^{-\eta/2}\right)$ based on the settings of $E_{1}$ and $E_{2}$. The regret of the leader can be bounded as:
@@
where (A) follows from Lemma C.2 and (B) uses the fact that $c\geq 1$ and $d\leq 1$. The regret of the follower can similarly be bounded as:
@@
where (B) follows from Lemma C.2. This proves the desired result.
E.2 Proof of Theorem 6.3
@@
The proof follows a similar argument to the proof of Theorem 4.4 and borrows some lemmas from Appendix C.3
@@
Assume that the clean event $G$ holds. This occurs with probability at least $1-(1+|\mathcal{A}|)T^{-3}$ (Lemma C.3), so the clean event not occurring counts negligibly towards regret.
@@
$\beta_{1}^{\text{tol}}\cdot(T-E\cdot\left|\mathcal{A}\right|)-\sum_{t=E\cdot|\mathcal{A}|}^{T}v_{1}(a_{t},b_{t})$
@@
$=\sum_{a\in\mathcal{A}}\sum_{t\in S(a)}\left(c\cdot(\epsilon^{*})^{d}+\max_{a\in\mathcal{A}}\min_{b\in\mathcal{B}_{\epsilon^{*}}(a)}v_{1}(a,b)-v_{1}(a_{t},b_{t})\right)$
$\leq|\mathcal{A}|+\sum_{a\in\mathcal{A}}\sum_{t\in S(a)\setminus\left\{\max(S(a))\right\}}\left(c\cdot(\epsilon^{*})^{d}+\max_{a\in\mathcal{A}}\min_{b\in\mathcal{B}_{\epsilon^{*}}(a)}v_{1}(a,b)-v_{1}(a_{t},b_{t})\right)$
@@
$+\underbrace{\sum_{a\in\mathcal{A}}(n_{E\cdot|\mathcal{A}|,T+1}(a)-1)\cdot\left(\max_{a\in\mathcal{A}}\min_{b\in\mathcal{B}_{\epsilon^{*}}(a)}v_{1}(a,b)-\frac{1}{n_{E\cdot|\mathcal{A}|,T+1}(a)-1}\sum_{t\in S(a)\setminus\left\{\max(S(a))\right\}}\left(v_{1}(a_{t},b_{t})\right)\right)}_{(2)}$
@@
$\sum_{a\in\mathcal{A}}(n_{E\cdot|\mathcal{A}|,T+1}(a)-1)\cdot\left(\max_{a\in\mathcal{A}}\min_{b\in\mathcal{B}_{\epsilon^{*}}(a)}v_{1}(a,b)-\frac{1}{n_{E\cdot|\mathcal{A}|,T+1}(a)-1}\sum_{t\in S(a)\setminus\left\{\max(S(a))\right\}}\left(v_{1}(a_{t},b_{t})\right)\right)$
$\leq\sum_{a\in\mathcal{A}}(n_{E\cdot|\mathcal{A}|,T+1}(a)-1)\cdot\frac{20\sqrt{\log T}}{\sqrt{n_{E\cdot|\mathcal{A}|,T+1}(a)}-1}$
$\leq O\left(\sqrt{\log T}\cdot\sum_{a\in\mathcal{A}}\sqrt{n_{E\cdot|\mathcal{A}|,T+1}(a)-1}\right)$
@@
where the first inequality uses Lemma C.4 and the last inequality uses Jensen’s inequality.
@@
However, by Lemma C.4, we can also lower bound the average reward across all time steps (except for the last time step) where $a^{\prime}$ is pulled in terms of $n_{E\cdot|\mathcal{A}|,T}(a^{\prime})$ as follows:
@@
Appendix F Proofs for Section 7
F.1 Proofs for Section 7.1
@@ Appendix F Proofs for Section 7
We use the following notation in the proof. Let $n_{t}(a)$ be the number of times that arm $a$ has been pulled up prior to time step $t$. Following Section A.4, the follower’s history can be represented as:
@@
$\mathbb{P}\left[\forall t\in[T],a\in\mathcal{A}\mid v_{2}(a_{t},b_{t})\geq\max_{b\in\mathcal{B}}v_{2}(a_{t},b)-g(n_{t+1}(a),T)\right]\geq 1-|\mathcal{A}|\cdot T^{-3},$
@@
Using Lemma F.1, it suffices to analyze the high-probability instantaneous regret and high-probability anytime regret of the following standard bandit algorithms as single-bandit learners with arms $\mathcal{B}$, mean rewards $v_{2}(b)$, and stochastic rewards $r_{2,t}(b)$. In the proofs, we let $n_{t}(b)$ denote the number of times that arm $b$ has been pulled prior to time step $t$.
@@
By Lemma F.1, it suffices to show the bound for ActiveArmElimination (using phase lengths $M_{i}=\Theta(\log(T)\cdot 2^{2i})$) as a single-bandit learner with arms $\mathcal{B}$, mean rewards $v_{2}(b)$, and stochastic rewards $r_{2,t}(b)$. We let $n_{t}(b)$ denote the number of times that arm $b$ has been pulled prior to time step $t$ in the current phase. Let $\hat{v}_{2,t}(b)$ denote the empirical mean reward for arm $b$ over the rewards observed prior to time step $t$ in the previous (last completed) phase. Let $\mathcal{B}^{\prime}_{t,\text{curr}}$ be the set of arms active in the current phase, and let $\mathcal{B}^{\prime}_{t,\text{prev}}$ be the set of arms active in the previous (last completed) phase. For each time step $t$, let $s^{\prime}_{t}$ denote the index of the previous (last completed) phase at time step $t$.
@@
By Lemma F.1, it suffices to show the instantaneous regret bound for ExploreThenCommit as a single-bandit learner with arms $\mathcal{B}$, mean rewards $v_{2}(b)$, and stochastic rewards $r_{2,t}(b)$. We let $n_{t}(b)$ denote the number of times that arm $b$ has been pulled prior to time step $t$. Let $\hat{v}_{2,t}(b)$ denote the empirical mean reward for arm $b$ over the rewards observed prior to time step $t$.
@@
By Lemma F.1, it suffices to show the anytime regret bound for UCB as a single-bandit learner with arms $\mathcal{B}$, mean rewards $v_{2}(b)$, and stochastic rewards $r_{2,t}(b)$. We let $n_{t}(b)$ denote the number of times that arm $b$ has been pulled prior to time step $t$. Let $\hat{v}_{2,t}(b)$ denote the empirical mean reward for arm $b$ over the rewards observed prior to time step $t$.
@@
We condition on $G$ for the remainder of the analysis. Since the arm with highest upper confidence bound is always chosen and since $G$ holds, the selected arm $b_{t}$’s true mean $v_{2}(b_{t})$ falls within the $2\cdot\sqrt{\frac{\log(T)}{n_{t}(b_{t})}}$ bound. By Lemma A.1, this means that the regret at any time step $t$ for any arm $a\in\mathcal{A}$ is upper bounded by: